OpenStudy (anonymous):

A ball is thrown into the air with an upward velocity of 20 feet per second. Its height, h, in feet after t seconds is given by the function h(t)=-16t^2+20t+2. What is the ball's maximus height? How long does it take the ball to reach its maximum height? Round to the nearest hundredth, if necessary. A. Reaches a max height of 8.25 feet after 1.25 seconds B. Reaches a max height of 8.25 feet after 0.63 seconds C. Reaches a max height of 0.16 feet after 1.34 seconds D. Reaches a max height of 0.32 feet after 1.34 seconds. I will give medal thanks!

OpenStudy (anonymous):

What is the vertex of the parabola \[ h(t)=-16t^2+20t+2. \]

OpenStudy (anonymous):

Or if you know derivative, what is the derivative of the above function?

OpenStudy (anonymous):

Look at the graph

OpenStudy (anonymous):

Without Calculus: We can write \[h(t)=2-16 \left(t^2-\frac{5 t}{4}+\frac{25}{64}-\frac{2 5}{64}\right)=\\-16 \left(t^2-\frac{5 t}{4}+\frac{25}{64}\right)+ \frac{25}{4}+2=\\ -16 \left(t-\frac{5}{8}\right )^2+\frac{33}{4} \]

OpenStudy (anonymous):

So, it attains it maximum height at t=5/8=.625 second and the maximum height is 33/4=8.25 ft

OpenStudy (anonymous):

well that jerk @-_llama_- never gave the medal he promised..... this helped me out so thank you @eliassaab